OpenStudy (anonymous):

is it true that fº(g+h) = (fºg)+(fºh). fºg is, for example f(g(x). Try, f(x)=sin x, g(x) = x, h(x) =x.

myininaya (myininaya):

f(g+h)=f(x+x)=sin(x+x)=sinxcosx+sinxcosx=2sinxcosx=sin(2x) f(g)+f(h)=sin(x)+sin(x)=2sin(x) does not equal sin(2x)

OpenStudy (anonymous):

but are you sure, by fºg I dont mean, f(x)*g(x), I mean f(g(x)) ??

myininaya (myininaya):

yes f(g)

myininaya (myininaya):

i write a function of a function not a function times a function

OpenStudy (anonymous):

ok